Half of a herd of deer are grazing in the field and three fourths of the remaining are playing nearby. The rest 9 are drinking water from the pond. Find the number of deer in the herd. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the total number of deer be x.
Deer grazing in the field = x/2
Deer playing nearby = x/2 × 3⁄4 = 3x/8
Deer drinking water = 9
According to the question,
x/2+3x/8+9=x
(4x+3x)/8+9=x
⇒7x/8+9=x
⇒x–7x/8=9
⇒(8x–7x)/8=9
⇒x=9×8
⇒x=72

Thanks